OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

@conner143 , @edwards555

OpenStudy (anonymous):

ok ill help

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

no problem anytime

OpenStudy (anonymous):

what u want

OpenStudy (anonymous):

got any ideas?

OpenStudy (anonymous):

Hi @FEARLESS_RAIN! Do you still need help?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Okay! So the first inequality says that y is LESS than or equal to x - 5. What we do to solve this is substitute the x value in each pair for the x in the inequality. Let's try the first pair, -5,2 where -5 is x. And we will see if the answer is less than or equal to the second number in the pair, which is 2. \[-5 - 5 = ?\] Can you tell me what that inequality is equal to?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

actually it is -10. Is -10 less than or equal to 2?

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@agent0smith @AMYCARTER

OpenStudy (anonymous):

You're right! Now let's try it for the next inequality, -x - 4. We substitute the same value we used for the other one, which is -5 and let's see if we get a value less than or equal to 2. \[-5 - 4 = ?\] What does that equal to?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wait a minute, I didn't read that inequality very well it says y is GREATER than or equal to -x - 4! I made a mistake, our answer can't be -5,2. We're going to have to move on to the next one, -5,-2

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

Or no let's do 5,-2. \[5 - 5 = 0\] obviously so 5,-2 can't be right either

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Lets try the next one then.

OpenStudy (anonymous):

The next one is -5,-2 so I'm going to plug in -5 for x which we already know comes out to -10 and -10 is LESS than -2 so we can do the next one, -x - 4. Let's see if y is greater than or less than this inequality. \[-5 - 4 = ?\]

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

Hmmm okay let's see if the next one is 5,2. 5-5 = 0 so that first part is right. -5 - 4 = -9 but that isn't right either....oh gosh this is confusing...

OpenStudy (anonymous):

Yeah ik

OpenStudy (anonymous):

I have no idea what it is

OpenStudy (anonymous):

I'm not sure either....why don't you just go with -5,-2? I don't guarantee it is the right answer though...I'm sooo sorry